Jump to main content.


Drake Chemical

EPA Daily Situation Reports for May 1998

U. S. ARMY CORPS OF ENGINEERS
DRAKE CHEMICAL SUPERFUND SITE

182 MYRTLE STREET
LOCK HAVEN, PA 17745
PHONE 717-748-4505
FAX 717-748-7549

Sunday, May 31, 1998

FOR DATES: From 7:00 a.m. Saturday, May 30, 1998 to 7:00 a.m. Sunday, May 31, 1998.

WEATHER: Partly cloudy. The temperature range was 62o F to 83o F.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 999.3 tons of soil was treated during the time frame of this report. Off-line time was attributed to weigh belt calibration, tripping of feed screws, and for maintenance of an excavator and power screen in the feed preparation building. Site soil is being processed at approximately 42 tons per hour. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued feed preparation, debris separation, ash handling, and ash backfilling. Continued off-site and perimeter air monitoring.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None

PREPARED BY: Rick Lupold, (USACE).


Saturday, May 30, 1998

FOR DATES: From 7:00 a.m. Friday, May 29, 1998 to 7:00 a.m. Saturday, May 30, 1998.

WEATHER: Clear in morning. Brief period of thunderstorm with high winds in late afternoon. The temperature range was 60o F to 89o F.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 19.9 hours and 822 tons of soil was treated during the time frame of this report. Off-line time was attributed to instrument calibration, monthly AWFSO interlock checks and shutdown of feed during a period of high wind with the possibility of lightning. Site soil is being processed at approximately 42 tons per hour. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued feed preparation, excavation, drum handling, debris separation, ash handling, ash backfilling and WWTP operation. Placed odor suppressant foam at excavation face and ConCover on feed material stockpile. Continued off-site and perimeter air monitoring. Received analytical for bot-ash-05-24-98-118, which passed the criteria for backfilling.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Kevin Fanning of CCDES.

PREPARED BY: Danny Zion for David Modricker (USACE).


Friday, May 29, 1998

FOR DATES: From 7:00 a.m. Thursday, May 28, 1998 to 7:00 a.m. Friday, May 29, 1998.

WEATHER: Clear. The temperature range was 55o F to 83o F.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.7 hours and 996.6 tons of soil was treated during the time frame of this report. Off-line time was attributed to weigh belt calibration and kiln pressure fluctuations. Site soil is being processed at approximately 42 tons per hour. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued feed preparation, excavation, drum handling, debris separation, ash handling, ash backfilling and WWTP operation. Placed odor suppressant foam at excavation face and ConCover on feed material stockpile. Continued off-site and perimeter air monitoring. Received analytical for bot-ash-05-23-98-117, which passed the criteria for backfilling.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandra Downs of PADEP, Kurt Davis of GAP.

PREPARED BY: David Modricker (USACE).


Thursday, May 28, 1998

FOR DATES: From 7:00 a.m. Wednesday, May 27, 1998 to 7:00 a.m. Thursday, May 28, 1998.

WEATHER: Clear. The temperature range was 51o F to 80o F.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 21.97 hours and 899 tons of soil was treated during the time frame of this report. Off-line time was attributed to weigh belt calibration, adjustment of the feed building lower weigh belt tensioner and kiln pressure spikes. Site soil is being processed at approximately 41 tons per hour. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued feed preparation, excavation, drum handling, debris separation, ash handling, ash backfilling and WWTP operation. Placed odor suppressant foam at excavation face. Continued off-site and perimeter air monitoring. Received analytical for bot-ash-05-23-98-116, which passed the criteria for backfilling.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Alan Roy and Lazon Allen of USACE.

PREPARED BY: David Modricker (USACE).


Wednesday, May 27, 1998

FOR DATES: From 7:00 a.m. Tuesday, May 26, 1998 to 7:00 a.m. Wednesday, May 27, 1998.

WEATHER: Clear. The temperature range was 51o F to 78o F.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 22.7 hours and 936.5 tons of soil were treated during the time frame of this report. Off-line time was attributed to weigh belt calibration, flameout of a secondary burner, and system pressure fluctuations. Site soil is being processed at approximately 41 tons per hour. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued feed preparation, excavation, drum handling, debris separation, ash handling, ash backfilling and WWTP operation. Placed concover on stockpile and placed odor suppressant foam at excavation face. Continued off-site and perimeter air monitoring. Received analytical for bot-ash-05-20-98-115 and fly-ash-05-20-98. Both samples passed the criteria for backfilling.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: S. Downs of PADEP, K. Davis of GAP and WTAM TV Newswatch reporter and cameraman.

PREPARED BY: David Modricker (USACE).


Tuesday, May 26, 1998

FOR DATES: From 7:00 a.m. Monday, May 25, 1998 to 7:00 a.m. Tuesday, May 26, 1998.

WEATHER: Rain in morning, Overcast. The temperature range was 53o F to 74o F.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.95 hours and 984.2 tons of soil were treated during the time frame of this report. Off-line time was attributed to weigh belt calibration. Site soil is being processed at approximately 41 tons per hour. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued feed preparation, debris separation and ash backfilling. Continued excavation and grading at ash backfill truck turnaround. Continued off-site and perimeter air monitoring.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: David Modricker (USACE).


Monday, May 25, 1998

FOR DATES: From 7:00 a.m. Sunday, May 24, 1998 to 7:00 a.m. Monday, May 25, 1998.

WEATHER: Cloudy with rain. The temperature range was 43o F to 76o F.

I. CURRENT SITUATION: The contractor continues to work 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on line for 23 hours, 0 minutes and processed approximately 964.1 tons of material. Site soils were being processed at approximately 41 tons per hour. Waste feed was stopped for a period of time due to thunder storms in the area. Offline time was also attributed to conducting a weekly weigh belt calibration. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations, feed preparation, debris separation, ash handling, and rock backfilling. Continued dust control operations by applying water to soils or applying Concover to the stockpile. Continued perimeter air monitoring. Received lab reports BOT-ASH-05-19-98-114 and FLY-ASH-05-16-98-111. The lab reports indicate that treated materials are within project specifications for backfilling.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, drum shipment off-site, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: John Ricard and David Modricker (USACE).


Sunday, May 24, 1998

FOR DATES: From 7:00 a.m. Saturday, May 23, 1998 to 7:00 a.m. Sunday, May 24, 1998.

WEATHER: Clear. The temperature range was 41o F to 75o F.

I. CURRENT SITUATION: The contractor continues to work 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on line for 21 hours, 28 minutes and processed approximately 809.9 tons of material. Site soils were being processed at approximately 38 tons per hour. Off line time is attributed to a PP&L power failure that interrupted power to the site at 2135 hrs. The temporary interruption in power caused an unplanned TRV opening that lasted approximately 45 seconds. All required incident notifications were made. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations, continued rock excavation from the truck turn around near the utility corridor for use as clean fill material in the backfill area above elevation 545.0., continued feed preparation, debris separation, ash handling, ash backfilling, and rock stockpiling. Continued with dust control operations by applying water to soils or applying Concover to the stockpile.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, drum shipment off-site, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: An unplanned TRV opening occurred as noted in section I above.

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson and David Modricker (USACE).


Saturday, May 23, 1998

FOR DATES: From 7:00 a.m. Friday, May 22, 1998 to 7:00 a.m. Saturday, May 23, 1998.

WEATHER: Partly cloudy. The temperature range was 39o F to 67o F.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) resumed contaminated soil feed at 0547 hours. The TDF was on line for 51 minutes and processed approximately 29 tons of material during the period of this report. The contractor continues to work 24 hours a day, seven days a week operating the TDF. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Finished TDF repairs and heatup to operating temperature. Began excavating clean rock fill material from the truck turn around area along the utility corridor for use as clean backfill material on site above elevation 545.0. Continued feed preparation, debris separation and ash handling activities. Continued the placement of odor suppressant foam within the exclusion zone and continued with dust control operations. Continued off-site and perimeter air monitoring programs. Received ash analytical report for BOT-ASH-05-17-98-112 that met site backfill requirements. Received ahs analytical report for BOT-ASH-05-18-98-113 (ash bin 5) which did not meet site specifications for backfill, and the material from ash bin 5 was placed back in the exclusion zone for further treatment. All other ash analytical tests during the operations phase have met project requirements for backfilling.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, drum shipment off-site, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Kurt Davis (GAP)

PREPARED BY: Steve Gibson and David Modricker (USACE).


Friday, May 22, 1998

FOR DATES: From 7:00 a.m. Thursday, May 21, 1998 to 7:00 a.m. Friday, May 22, 1998.

WEATHER: Clear. The temperature range was 46o F to 75o F.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) is currently in a cold shut down to perform maintenance on the acid gas scrubber system. A fiberglass header pipe in the scrubber tank is being repaired. The contractor continues to work 24 hours a day, seven days a week maintaining the TDF. Related ongoing activities include soil feed preparation, excavation, and ash handling.

II. ACTIONS TAKEN: Continued TDF repairs, materials excavation, feed preparation, debris separation, ash handling, and ash backfilling. Continued the placement of odor suppressant foam within the exclusion zone and continued with dust control operations. Continued off-site and perimeter air monitoring. Performed some minor undercutting below elevation 545.0 within the exclusion zone which revealed no additional visibly contaminated soil.

III. FUTURE PLANS: Expect to begin TDF heat up and resume waste feed to the TDF on Friday, May 22, 1998. Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, drum shipment off-site, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandy Downs (PADEP), Chuck Sweeny (Rotary Club), Mustafa O. (Rotary International)

PREPARED BY: Steve Gibson and David Modricker (USACE).


Thursday, May 21, 1998

FOR DATES: From 7:00 a.m. Wednesday, May 20, 1998 to 7:00 a.m. Thursday, May 21, 1998.

WEATHER: Clear to partly cloudy. The temperature range was 61o F to 88o F.

I. CURRENT SITUATION: As of Thursday morning, May 21, 1998, the Thermal Destruction Facility (TDF) is currently off line to perform repairs to the acid gas scrubber system piping. The contractor continues to work 24 hours a day, seven days a week operating the TDF. The TDF was on-line for 7.2 hours and 307.9 tons of soil were treated at approximately 43 tons per hour during the time frame of this report. Off-line time was attributed to repairs to the acid gas scrubber system. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued materials excavation, feed preparation, debris separation, ash handling, and ash backfilling. Continued the placement of odor suppressant foam within the exclusion zone and continued with dust control operations. Continued off-site and perimeter air monitoring.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, drum shipment off-site, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: The contractor's control room operators observed unusually high flow rates in one of the two acid gas steam scrubbers during operations. Feed was stopped at 1409 hours and the Thermal Relief Vent (TRV) was opened at 1522 hours, once the proper authorities had been notified. The high flow rates within the scrubber system were the result of a crack in the pipe header at the flange attachment to the scrubber tank. The cracked pipe is scheduled to be repaired on Thursday, May 21, 1998. TDF operations are expected to resume on Friday, May 22, 1998.

V. NOTED VISITORS: Kurt Davis (GAP)

PREPARED BY: Steve Gibson and David Modricker (USACE).


Wednesday, May 20, 1998

FOR DATES: From 7:00 a.m. Tuesday, May 19, 1998 to 7:00 a.m. Wednesday, May 20, 1998.

WEATHER: Clear. The temperature range was 58o F to 89o F.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). Related ongoing activities include soil feed preparation and ash handling. The TDF was on-line for 23.4 hours and 1015.2 tons of soil were treated during the time frame of this report. Off-line time was attributed to the daily weight belt calibration. Site soil is being processed at approximately 43 tons per hour.

II. ACTIONS TAKEN: Continued TDF operations. Continued materials excavation, feed preparation, debris separation, ash handling, and ash backfilling. Continued the placement of odor suppressant foam within the exclusion zone and continued with dust control operations. Some dust was observed blowing across the Drake site from off site. Appropriate notifications were made and action taken. Continued off-site and perimeter air monitoring. Received analysis for bottom ash sample bot-ash-5-15-98-111. The ash sample met the project requirements and specifications for backfilling.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Mike Welch, Larry Newcomer, Steven Webster, Sandy Downs (PADEP)

PREPARED BY: Steve Gibson and David Modricker (USACE).


Tuesday, May 19, 1998

FOR DATES: From 7:00 a.m. Monday, May 18, 1998 to 7:00 a.m. Tuesday, May 19, 1998.

WEATHER: Clear. The temperature range was 53o F to 85o F.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). Related ongoing activities include soil feed preparation and ash handling. The TDF was on-line for 22.7 hours and 996.5 tons of soil were treated during the time frame of this report. Off-line time was attributed to weight belt calibrations, low combustion efficiency, and chiller repair for the continuous emissions monitors (CEMs). Site soil is being processed at approximately 44 tons per hour.

II. ACTIONS TAKEN: Continued TDF operations. Continued feed preparation, debris separation and ash backfilling. Continued off-site and perimeter air monitoring. Received analysis for bottom ash sample bot-ash-5-14-98-110. The ash sample met the project requirements and specifications for backfilling.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination

IV. PROBLEMS NOTED OR ANTICIPATED: Feed was stopped at 1900 hours to repair the chiller system that cools the CEM sample stack gases. Feed was restarted one hour later once repairs had been made and the CEMs were back on-line.

V. NOTED VISITORS: Dick Lee (USACE), Sandy Downs (PADEP), Kurt Davis (GAP)

PREPARED BY: Steve Gibson and David Modricker (USACE).


Monday, May 18, 1998

FOR DATES: From 7:00 a.m. Sunday, May 17, 1998 to 7:00 a.m. Monday, May 18, 1998.

WEATHER: Clear. The temperature range was 50o F to 84o F.

I. CURRENT SITUATION: Waste feed resumed at 2034 hours on Sunday after repairs were made to the acid gas scrubber system. The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). Related ongoing activities include soil feed preparation and ash handling. The TDF was on-line for 10.43 hours and treated 419.5 tons of soil during the time frame of this report. Off-line time was attributed to replacement of the acid gas scrubber pump. Site soil is being processed at approximately 40 tons per hour.

II. ACTIONS TAKEN: Continued TDF operations. Continued feed preparation, debris separation and ash backfilling. Continued off-site and perimeter air monitoring. Received analysis for bottom ash samples 05-11-98-108 and 5-13-98-109. Both samples met requirements for backfilling.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling and backfilling, drum handling, off-site and perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: David Modricker (USACE).


Sunday, May 17, 1998

FOR DATES: From 7:00 a.m. Saturday, May 16, 1998 to 7:00 a.m. Sunday, May 17, 1998.

WEATHER: Clear and Sunny. The temperature range was 50.4( F to 86( F.

I. CURRENT SITUATION: Waste feed to the Thermal Destruction Facility (TDF) was suspended for the 24 hour period of this report while maintenance and repairs were made to the acid gas scrubber system.

II. ACTIONS TAKEN: Continued removal of drum carcasses from the excavation area within the exclusion zone. Placed stone and rubble at the bottom of the recently excavated area and continued ash backfilling procedures. Placed concover on excavated soil stockpile material. Continued soil screening, debris separation, rock washing, and inorganic debris decontamination activities. Continued WWTP operations. Continued off-site and perimeter air monitoring programs.

Repairs were made to one of the acid gas scrubber pumps. Inspection and maintenence of the acid gas scrubber system continued. The Thermal Relief Valve remained open until 0557 hours, when the kiln burners and secondary combustion chamber burners were turned back on to heat up the TDF in preparation to resume operations.

III. FUTURE PLANS: Major future activities include: bringing the TDF back on line and continuing TDF operations; continuation of excavation, feed preparation, debris handling, ash handling & backfilling, drum handling, off-site and perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: David Modricker (USACE).


Saturday, May 16, 1998

FOR DATES: From 7:00 a.m. Friday, May 15, 1998 to 7:00 a.m. Saturday, May 16, 1998.

WEATHER: Clear and Sunny. The temperature range was 52* F to 85* F.

I. CURRENT SITUATION: The contractor continues to work 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF) and performing related activities such as soil feed preparation and ash handling. The TDF was on-line for 7.4 hours and processed approximately 299 tons of material during the time period covered under this report. Site soil is being processed at a rate of approximately 40 tons per hour. Off-line time is attributed to maintenance on the acid gas scrubber system.

II. ACTIONS TAKEN: Continued processing site soils through the TDF. Continued excavating contaminated soils in the exclusion zone. Continued the excavation of drums and drum carcasses. Continued overpacking, sampling, and labeling drums for shipment off-site. Placed concover on excavated soil stockpile material, and placed odor suppressant foam over excavation faces. Continued soil screening, debris separation, rock washing, and inorganic debris decontamination activities. Continued WWTP operations. Continued off-site and perimeter air monitoring programs.

Incinerator operators noticed increased temperatures in the acid gas scrubber system on Friday morning. To cool the acid gas scrubber system and perform inspections, waste feed was stopped at 1046 hours, and the TRV was opened at 1243 hours. Appropriate notifications were made for the planned TRV opening. It was determined that rubber material had clogged some of the venturri spray nozzels that cool the gasses in the acid gass scrubbers. The nozzles were cleaned, the TDF was brought up to operating temperatures, and waste feed resumed at 1900 hours.

During the evening, incinerator operators again noticed increased temperatures in the acid gas scrubber system. To cool the acid gas scrubber system and perform inspections, waste feed was stopped at 2239 hours, and the TRV was opened at 2356 hours. Appropriate notifications were made for the planned TRV opening. It was again determined that rubber material had clogged some of the venturri spray nozzels that cool the gasses in the acid gass scrubbers. A more detailed investigation was conducted on Saturday to determine the source of the rubber material. It was later determined that the rubber material came from the rubber lined housing of the acid gas scrubber pumps. The scrubber system was drained and the contractor cleaned and flushed the system. Waste feed remained off for the remainder of this reporting period.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, feed preparation, debris handling, ash handling & backfilling, drum handling, off-site and perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: Waste feed operations were suspended for a period of time as noted above to perform maintenence and repairs to the acid gas scrubber system. To facilitate the repairs, two planned Thermal Relief Valve (TRV) openings ocurred.

V. NOTED VISITORS: Jim Cheney (USACE), Mike Welch (PADEP), Sandy Downs (PADEP), Kurt Davis (GAP).

PREPARED BY: David Modricker (USACE).


Friday, May 15, 1998

FOR DATES: From 7:00 a.m. Thursday, May 14, 1998 to 7:00 a.m. Friday, May 15, 1998.

WEATHER: Clear and Sunny. The temperature range was 45o F to 81o F.

I. CURRENT SITUATION: The contractor continues to work 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF) and performing related activities such as soil feed preparation and ash handling. The TDF was on-line for 23.03 hours and processed approximately 962.7 tons of material during the time period covered under this report. Site soil is being processed at a rate of approximately 42 tons per hour. Off-line time is attributed to weight belt calibrations, and kiln burner maintenance. Stack dioxin testing is being performed throughout the time frame of this report.

II. ACTIONS TAKEN: Continued processing site soils through the TDF. Received lab reports for BOT-ASH-05-09-98-106A and BOT-ASH-05-10-98-107. The lab reports indicate that treated materials are within project specifications for backfilling. Continued excavating contaminated soils in the exclusion zone. Continued the excavation of drums and drum carcasses. Continued overpacking, sampling, and labeling drums for shipment off-site. Placed concover on excavated soil stockpile material. Continued soil screening, debris separation, rock washing, and inorganic debris decontamination activities. Continued WWTP operations. Continued off-site and perimeter air monitoring programs.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, feed preparation, debris handling, ash handling & backfilling, drum handling, off-site and perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Jim Cheney (USACE), Sandy Downs (PADEP), Kurt Davis (GAP).

PREPARED BY: David Modricker (USACE).


Thursday, May 14, 1998

FOR DATES: From 7:00 a.m. Wednesday, May 13, 1998 to 7:00 a.m. Thursday, May 14, 1998.

WEATHER: Overcast, clear in afternoon. The temperature range was 42o F to 74o F.

I. CURRENT SITUATION: The contractor continues to work 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF) and performing related activities such as soil feed preparation and ash handling. The TDF was on-line for 14.48 hours and processed approximately 588.1 tons of material during the period covered under this report. Site soil is being processed at a rate of approximately 41 tons per hour. Off-line time is attributed to weight belt calibrations, the completion of the feed screw maintenance, and feed screw adjustments.

II. ACTIONS TAKEN: Continued processing site soils through the TDF. Continued backfilling of ash that met project backfill requirements. Continued excavating contaminated soils in the exclusion zone. Continued the excavation of drums and drum carcasses. Continued overpacking, sampling, and labeling drums for shipment off-site. Placed foam odor suppressant over recently excavated areas. Continued soil screening, debris separation, and rock washing & inorganic debris decontamination activities. Continued off-site and perimeter air monitoring programs.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, feed preparation, debris handling, ash handling and backfilling, drum handling, off-site and perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination. Stack testing is scheduled to begin 0730 hours, Thursday, May 14, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Jim Cheney of USACE, Kurt Davis of GAP.

PREPARED BY: David Modricker (USACE).


Wednesday, May 13, 1998

FOR DATES: From 7:00 a.m. Tuesday, May 12, 1998 to 7:00 a.m. Wednesday, May 13, 1998.

WEATHER: Overcast and rain. The temperature range was 47o F to 63o F.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) is currently off-line to allow for replacement of the kiln feed screws and the kiln feed screw water cooling jacket housing. The contractor continues to work 24 hours a day, seven days a week operating the TDF and performing related activities such as soil feed preparation and ash handling. TDF operations are expected to resume today.

II. ACTIONS TAKEN: The TDF was off-line during the time frame of this report for replacement of the kiln feed screws and the kiln feed screw cooling jacket. Received analysis for bottom ash sample 05-08-98-105 which met requirements for backfilling. Backfilling ash that met backfill requirements. Excavating contaminated soils in the exclusion zone. Placed foam odor suppressant. Continued soil screening, debris separation, and rock washing & inorganic debris decontamination activities. Continued off-site and perimeter air monitoring programs

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, feed preparation, debris handling, ash handling and backfilling, drum handling, off-site and perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination. Stack testing is scheduled to begin the evening of Wednesday, May 13, 1998 or Thursday, May 14, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: A crack was found in the kiln feed screw cooling jacket during water pressure testing. Replacement of the jacket is in process.

V. NOTED VISITORS: Jim Cheney of USACE, Kurt Davis of GAP.

PREPARED BY: David Modricker (USACE).


Tuesday, May 12, 1998

No report available.


Monday, May 11, 1998

No report available.


Sunday, May 10, 1998

No report available.


Saturday, May 9, 1998

No report available.


Friday, May 8, 1998

No report available.


Thursday, May 7, 1998

No report available.


Wednesday, May 6, 1998

No report available.


Tuesday, May 5, 1998

No report available.


Monday, May 4, 1998

No report available.


Sunday, May 3, 1998

No report available.


Saturday, May 2, 1998

No report available.


Friday, May 1, 1998

No report available.

Region 3 | Mid-Atlantic Cleanup | Mid-Atlantic Superfund |EPA Home | EPA Superfund Homepage


Local Navigation


Jump to main content.